OpenStudy (anonymous):

solve the equation. 16=4x-8

OpenStudy (anonymous):

Add 8 to both sides.

OpenStudy (anonymous):

Add 8 to both sides then divide by 4

OpenStudy (anonymous):

24=4x, from here, divide by 4/ x=24/4=6, x=6

OpenStudy (anonymous):

16=4x-8 add 8 to both sides to cancel out the -8 24=4x x=6

OpenStudy (anonymous):

\[16=4x-8\]Add 8 to both sides.\[24=4x\]Divide both sides by 4.\[6=x\]